OpenStudy (anonymous):

Can someone explain, Differentiation. the rule: U^n = nu^n-1 u`

sam (.sam.):

Looks like power rule

OpenStudy (anonymous):

I donow what questions were for homework. And nt know what questions were for homework. And nt know what questions were for homework. And no one telling me

OpenStudy (anonymous):

hi for example : (1+x^3)^5 ---> 5(3x^2)(1+x^3)^4

sam (.sam.):

\[\frac{d}{dx}(U^N)=NU^{N-1}(U')\] Applies to @Amirsd example

sam (.sam.):

U being \(1+x^3\), and N=5, U' is the differentiation of U, which you will get \(3x^2\)